A and b are two events such that p(a)=0.3 and p(AUB)=0.8 if a and b are two independent events then find p(b)

Grade:10

1 Answers

Nandana
7 years ago
Hi,
 Given ,
   p(A U B) = 0.8
   p(A) =0.3
   & A and B are independent events , so p(AB) = P(A)p(B)
   p(A U B) = p(A) + p(B) – p(AB)
    0.8 = p(A) + p(B) – p(A) p(B)
     0.8-0.3 = p(B) [1- 0.3]
from the above ,
   p(B) = 0.714285714 
 appx. p(B) = 0.714